OpenStudy (anonymous):

Ammonia is produced by the reaction of nitrogen gas with hydrogen gas. The balanced reaction is: N2 + 3H2 --> 2NH3. If you start with 19.64 g of H2, how many grams of NH3 can be produced?

OpenStudy (aaronq):

Convert the mass of reactant given to moles and use the stoichiometric coefficients to find moles product made. Set up a ratio using the species of interest, like so: e.g. for a general reaction: \(\color{red}{a}A + \color{blue}{b}B\) \(\rightleftharpoons\) \( \color{green}{c}C\) where upper case are the species (A,B,C), and lower case (a,b,c) are the coefficients , \(\dfrac{n_A}{\color{red}{a}}=\dfrac{n_B}{\color{blue}{b}}=\dfrac{n_C}{\color{green}{c}}\) From here you can isolate what you need. For example: if you have 2 moles of B, how many moles of C can you produce? solve algebraically: \(\dfrac{2}{\color{blue}{b}}=\dfrac{n_C}{\color{green}{c}}\rightarrow n_C=\dfrac{2*\color{green}{c}}{\color{blue}{b}}\) -------------------------------------------------------- To convert mass to moles, use the relationship: \(n=\dfrac{m}{M}\) where, M=molar mass, m=mass, and n= moles.

OpenStudy (anonymous):

How do I convert the reactants mass to moles and then use the equation for general reaction you provided.

OpenStudy (aaronq):

use the equation at the very bottom, it works to convert mass to moles or moles to mass.